Given the score X=72, σ=4 and z=1.64, the value of μ is

1
Expert's answer
2022-02-16T16:54:31-0500

we use z calculations as below

define z = ( x- u )/(σ)

substituting in the above formula we have

1.64 = ( 72 - u ) / ( 4 )

solving for u from the above we obtain that u = ( 72 - 6.56 ) = 65.44 which is the required solution.